Circular No. 8362

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004cv, 2004cw, 2004cx Three apparent supernovae have been discovered on unfiltered CCD

images — 2004cv by M. Armstrong (Rolvenden, U.K., 0.35-m reflector), 2004cw by C. Jacques (Belo Horizonte, Brazil), and 2004cx by L. A. G. Monard (Pretoria, South Africa, 0.30-m reflector).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004cv June 24.97 16^{h}06^{m}14.^{s}45 +18^{o}24′54′′.0 16.8 1′′ W, 2′′ N
2004cw June 13.24 19 30 57.45 −53 32 28.7 16.6 11′′.4 W, 8′′.3 N
2004cx June 26.06 23 47 52.86 −30 31 32.6 17.0 14′′ E, 17′′ S

Additional magnitudes by the respective discoverers: SN 2004cv in MCG +03-41-120, May 18 UT, [19.0; June 8, [19.0; 25.921, 16.6. SN 2004cw in ESO 184-75, 2002 Apr. 2.27, [17.8; 2003 Apr. 4.41, [18.7; June 19.18, 17.0; 25.03, 17.2. SN 2004cx in NGC 7755, May 29.1, [17.7; June 27.08, 16.8. Nothing is visible at the location of SN 2004cv on red Palomar Sky Survey images from 1987 and 1991 (limiting mag ∼ 20.0–20.8) and a blue POSS image from 1992. There appears to be a strong H II region at the location of SN 2004cw. Nothing is visible at the location of SN 2004cx on the Digitized Sky Survey (limiting red mag 20.5).

SUPERNOVAE 2003lx AND 2003ly Further to IAUC 8359, A. Connolly and the SDSS Collaboration report

two additional supernovae from last year.

SN  2003 UT  α_{2000}  δ_{2000}  r  Type (Age)
2003lx Apr. 24.44 16^{h}19^{m}21.^{s}66 +41^{o}05′23′′.6 18.5 Ia (+10)
2003ly Dec. 20.53 9 37 49.93 +10 11 38.1 20.4 Ia (+12)

Nothing was visible at the locations of the supernovae on earlier SDSS images taken on 2002 June 6 (2003lx) and 2003 Jan. 28 (2003lz).

COMET C/2002 T7 (LINEAR) Recent visual total-magnitude estimates: May 19.35 UT, 2.4 (D. A. J.

Seargent, Cowra, N.S.W., naked eye); June 4.47, 5.4 (Y. Nagai, Nagano, Japan, 12×50 binoculars); 15.90, 6.4 (J. J. Gonzalez, Leon, Spain, 7×50 binoculars); 20.34, 7.6 (Seargent, The Entrance, N.S.W., 10×50 binoculars); 24.91, 7.5 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars).

2004 June 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green